can i get help with solving the problem
-5+2LN3X=5

Answer by lwsshak3(11628)

can i get help with solving the problem
-5+2LN3X=5
..
-5+2ln3x=5
2ln3x=10
ln3x=5
convert to exponential form
e^5=3x
x=e^5/3=49.47
..
check:
-5+2ln(3*49.7)=4.9999...